Find the x- and y- intercepts of -2x+4y=8. Use these intercepts to graph the line.
Answer by rfer(16322)

First put in slope intercept form.
y=mx+b
4y=2x+8
y=2/4x+8/4
y=2/4x+2
--------------
slope m=2/4
y int b=2